# I2S-TFCKD: Intra-Inter Set Knowledge Distillation with Time-Frequency Calibration for Speech Enhancement

### v1

@@ III-C Time-frequency Cross-calibrated Multi-layer Matching Distillation
${\cal L}_{{\rm{TFCKD}}}$ $=\sum\limits_{{l_{s}}=1}^{{L_{s}}}\sum\limits_{{l_{t}}=1}^{{L_{t}}}\left(\alpha_{\left({{l_{s}},{l_{t}}}\right)}^{T}\left({{\bf{P}}_{{l_{t}}}^{T}-{\bf{P}}_{{l_{s}}}^{T}}\right)\log\left({\frac{{{\bf{P}}_{{l_{t}}}^{T}}}{{{\bf{P}}_{{l_{s}}}^{T}}}}\right)\right.$ (5)
$\left.+\alpha_{\left({{l_{s}},{l_{t}}}\right)}^{F}\left({{\bf{P}}_{{l_{t}}}^{F}-{\bf{P}}_{{l_{s}}}^{F}}\right)\log\left({\frac{{{\bf{P}}_{{l_{t}}}^{F}}}{{{\bf{P}}_{{l_{s}}}^{F}}}}\right)\right),$
@@
III-D Residual Fusion Features for Correlated Sets
${{{\bf{\tilde{R}}}}^{j-1}}={\rm{2DConv}}\left({{{\bf{R}}^{j-1}}}\right),$ (6)
${{{\bf{\tilde{F}}}}^{j}}={\rm{UpSampling}}\left({{{\bf{F}}^{j}}}\right),$
${{\bf{T}}^{j}}={\rm{2DConv}}\left({{{\bf{R}}^{j}}}\right).$ (9)
For each correlated set, we select the last recursion output as the representative fusion feature of the current set, which is then integrated to form a fused feature set. We believe that due to the symmetry of the encoder-decoder architecture in SE models, higher-level features from the middle layers exhibit stronger capability to learn effective information from lower-level features near the input/output ends. Therefore, in terms of residual fusion direction, the encoder and intermediate F-T processing modules follow a forward order starting from the first layer, while the decoder adopts a reverse order integrating from the final layer towards the middle. Corresponding representative features are generated for each correlated set in both the teacher and student models following the above process, and integrated into the teacher-student fused feature set for inter-set distillation.
III-E Training Procedure
In this section, we elaborate on the entire training procedure for intra-inter set distillation. The overall workflow is outlined in Algorithm 1. First, we pretrain the teacher network using the multi-resolution short-time Fourier transform (MRSTFT) loss as the backbone loss. The teacher is then frozen to serve as a guide for the student, and KD proceeds concurrently with the student’s training. Intra-set distillation is applied to the encoder, decoder, and F-T processing modules. Within each correlated set, besides calculating the intra-set time-frequency calibration distillation loss ${{\cal L}_{Intra}}$, we generate representative teacher-student fused features for the current set through layer-wise recursive computation. Inter-set distillation loss ${{\cal L}_{Inter}}$ is derived via cross-layer interactions within the fused feature set. The total loss for training the student model is formulated as:
${{\cal L}_{stu}}$ $={{\cal L}_{MRSTFT}}+\sum\limits_{k=1}^{q}{\sum\limits_{i=1}^{m}{\sum\limits_{j=1}^{n}{{{\cal L}_{Intra}}\left({{\bf{F}}_{s}^{i},{\bf{F}}_{t}^{j}}\right)}}}$ (10)
$+\sum\limits_{i=1}^{q}{\sum\limits_{j=1}^{q}{{{\cal L}_{Inter}}\left({{\bf{u}}_{s}^{i},{\bf{u}}_{t}^{j}}\right)}},$
IV Experimental Settings
IV-A Dataset Setup
@@
/* Iterate through all correlated sets. */
1 for $k\leftarrow 1$ to $q$ do
@@ 4       ${{\cal L}_{Intra}}=0$;
       /* Iterate through all teacher-student layers in the current set. */
5       for $i\leftarrow 1$ to $m$ do
@@ 7                   ${{\cal L}_{Intra}}+={\rm{TF\_Calibration}}\left({{\bf{F}}_{s}^{i},{\bf{F}}_{t}^{j}}\right)$;
8                   if $i==m$ then // Ensure the uniqueness.
@@ 29      ${\bf{u}}_{s}^{k}={\bf{T}}_{s}^{m}$;
       // The final output is chosen as the representative fused feature.
@@ 30       ${\bf{u}}_{t}^{k}={\bf{T}}_{t}^{n}$;
/* Iterate through fused features. */
34 for $i\leftarrow 1$ to $q$ do
@@ 36             ${{\cal L}_{Inter}}+={\rm{TF\_Calibration}}\left({{\bf{u}}_{s}^{i},{\bf{u}}_{t}^{j}}\right)$;
Algorithm 1 Intra-inter Set Distillation Processing
@@
IV-B Implementation Details
TABLE I: Baseline model hyperparameter settings.
Layer Teacher Student
@@ IV-B Implementation Details
Dilated Block (4$\times$×Conv2d) (2, 3), (1, 1), 128 $p$: [1, 2, 4, 8] (128, $T$, $F/4$) (128, $T$, $F/4$) $(2,3)$, $(1,1)$, 64 $p$: [1, 2, 4, 8] (64, $T$, $F/4$) (64, $T$, $F/4$)
Frequency-Time Block (128 GRU×$\times$2)$\times$4 (128, $T$, $F/4$) (128, $T$, $F/4$) 64 GRU×$\times$2 (64, $T$, $F/4$) (64, $T$, $F/4$)
Dilated Block (4$\times$×Conv2d) (2, 3), (1, 1), 128 $p$: [1, 2, 4, 8] (128, $T$, $F/4$) (128, $T$, $F/4$) $(2,3)$, $(1,1)$, 64 $p$: [1, 2, 4, 8] (64, $T$, $F/4$) (64, $T$, $F/4$)
@@
TABLE II: Distillation module hyperparameter settings.
@@
Layer Time Flow Frequency Flow
@@
Linear $T*factor$ units ($B$, $T$, $T$) ($B$, $T$, $T*factor$) $B*factor$ units ($T$, $B$, $B$) ($T$, $B$, $B*factor$)
ReLU - ($B$, $T$, $T*factor$) ($B$, $T$, $T*factor$) - ($T$, $B$, $B*factor$) ($T$, $B$, $B*factor$)
Linear $T$ units ($B$, $T$, $T*factor$) ($B$, $T$, $T$) $B$ units ($T$, $B$, $B*factor$) ($T$, $B$, $B$)
${l^{2}}$orm - ($B$, $T$, $T$) ($B$, $T$, $T$) - ($T$, $B$, $B$) ($T$, $B$, $B$)
@@
Layer Teacher Student
@@
In this paper, we use the DPDCRN model as our baseline, which takes complex spectrograms as input and outputs complex ratio masks. For the single-channel DNS dataset, the input channels are set to 2, while for the multi-channel L3DAS23 dataset, the input channels are set to 16. The hyperparameters of the teacher and student models are shown in Table I, where $C$ represents the number of input channels, $T$ is the number of speech frames, and $F$ is the frequency dimension of speech features. For the encoder and decoder parts, the number of channels in each convolutional layer of the teacher is twice that of the student. In Table I, $p$ represents the dilation factor, and both the teacher and student models use four types of dilation receptive fields. For the intermediate part, the teacher model employs four F-T modules, while the student model uses only one F-T module, with the number of GRU hidden units also being half those of the teacher model’s.
The hyperparameter settings for the trainable layers in the proposed distillation components are shown in Table II. The trainable components contain the computation of query and key embedding matrices in time-frequency calibration, as well as the generation process of recursively fused features for each correlated set. In the time-frequency calibration, the similarity embedding mappings are computed separately from two flow directions, with each flow containing two fully-connected layers for dimension transformation. The dimension transformation coefficient $factor$ in Table II is set to four in this paper. Residual fusion integrates the current layer features with the recursive features inherited from the previous layer via trainable parameters. In Table II, $C_{t/s}^{j}$ represents the number of teacher-student channels in the current layer, $F_{t/s}^{j}$ denotes the feature dimension of teacher-student features in the current layer, and $F_{t/s}^{j-1}$ denotes the feature dimension in the previous layer. The number of channels for recursive features of the teacher and student are set to 128 and 64, respectively. Notably, the trainable parameters of the distillation components are only involved in the training of the student model and do not impose additional burden during inference.
All speech signals are sampled at 16 kHz and segmented into 2.5-second chunks. For all models, the window length is set to 32 ms with a frame shift of 16 ms, and the FFT size is 512 points. We employ the MRSTFT loss [31] as the backbone loss. The models are implemented using PyTorch. The training configuration includes a learning rate of 0.0006 with the Adam optimizer, a batch size of 8, and a total training duration of 20 epochs. We provide source codes to ensure reproducibility [^3] The codes are available at https://github.com/JMCheng-SEU/I2S-TFCKD-SE..
IV-C Comparative Methods
@@
IV-D Evaluation Metrics
In this paper, we use six objective speech quality evaluation metrics, including the perceptual evaluation of speech quality (PESQ) [46], short-time objective intelligibility (STOI) [47], scale-invariant signal-to-noise ratio (SI-SNR) [48], and three mean opinion ratings on quality scales [49]: signal distortion (CSIG), the intrusiveness of background noise (CBAK), and overall quality (COVL). For all these metrics, higher scores indicate better performance. Additionally, on the L3DAS23 dataset, we use the official recommended word error rate (WER) computed via Wav2Vec [50] to evaluate SE effects for speech recognition purposes, where lower values indicate better performance.
V Experimental Results and Discussion
V-A Ablation Study
We conduct ablation experiments for the distillation components on both the validation set and the multi-SNR test set to verify the effectiveness of the proposed modules step-by-step. First, we set the M1 method as the basic form of feature distillation, adopting the teacher-student layer-wise distillation framework of FitNet. Then, method M2 is built upon the cross-layer attention-calibration distillation framework proposed by SemCKD, leveraging multi-layer teacher feature information to guide the student model. Method M3 expands on M2’s framework by introducing the multi-layer matching distillation with time-frequency calibration proposed in Section III-C of this paper. Finally, the M4 method (i. e., the proposed I2S-TFCKD framework) adds correlated set residual fusion to M3, and performs intra-inter set multi-layer distillation.
@@ V-A Ablation Study
V-A1 Analysis of Training Curves Trend
Fig. 4(a) shows the trend of each method’s backbone loss with the training epochs on the validation set. It can be observed that the feature distillation M1 enables the student’s backbone loss to decrease rapidly in the early training stages, but this initial advantage does not persist until model convergence. In contrast, method M2 achieves a lower converged backbone loss, indicating that cross-layer interactions provide additional knowledge to sustain the student’s learning. After introducing the time-frequency cross-calibration distillation loss in the M3 method, the variation of backbone loss with training epochs becomes more uniform and converges better, indicating that assigning multi-layer distillation weights based on time-frequency calibration leads to more stable knowledge transfer. On this basis, the M4 method introducing residual fusion distillation between correlated sets leads to lower convergence of the backbone loss. This demonstrates that intra-set and inter-set distillation complement each other, expanding the depth of knowledge perceived by the student model. Fig. 4(b) shows the variation of the speech metric PESQ with training epochs. The general performance order of different distillation methods is M1<M2<M3<M4, consistent with that in Fig. 4(a).
V-A2 Heatmap Visualization Analysis of Time-frequency Calibration Weights
V-A3 Analysis of Evaluation Metrics
TABLE III: Speech evaluation metrics for ablation comparisons of distillation components under different SNRs.
Methods Feature distillation Cross-layer distillation T-F calibration Intra-inter interactions Metric
@@ V-A3 Analysis of Evaluation Metrics
-M1 - - - 1.966 3.377 2.840 2.657 0.870 0.844
-M2 - - - 2.003 3.437 2.884 2.708 0.874 11.020
-M3 - - 2.057 3.500 2.894 2.767 0.880 11.107
-M4 - 2.106 3.567 2.950 2.830 0.886 11.253
@@
-M1 - - - 2.329 3.772 3.224 3.059 0.920 14.106
-M2 - - - 2.368 3.826 3.262 3.107 0.924 14.237
-M3 - - 2.426 3.904 3.274 3.174 0.928 14.270
-M4 - 2.491 3.954 3.333 3.238 0.931 14.461
@@
-M1 - - - 2.700 4.147 3.591 3.452 0.952 17.041
-M2 - - - 2.724 4.180 3.617 3.483 0.953 17.149
-M3 - - 2.788 4.244 3.633 3.546 0.956 17.157
-M4 - 2.850 4.285 3.686 3.606 0.958 17.356
@@
-M1 - - - 2.332 ($\ast$)1 3.765 ($\ast$) 3.218 ($\ast$) 3.056 ($\ast$) 0.914 ($\ast$) 13.997 ($\ast$)

### v2

@@ 3.3 Time-frequency cross calibration for multi-layer hierarchical matching
${\cal L}_{{\rm{TFCKD}}}=\sum\limits_{{l_{s}}=1}^{{L_{s}}}\sum\limits_{{l_{t}}=1}^{{L_{t}}}\left(\alpha_{\left({{l_{s}},{l_{t}}}\right)}^{T}\left({{\bf{P}}_{{l_{t}}}^{T}-{\bf{P}}_{{l_{s}}}^{T}}\right)\log\left({\frac{{{\bf{P}}_{{l_{t}}}^{T}}}{{{\bf{P}}_{{l_{s}}}^{T}}}}\right)+\alpha_{\left({{l_{s}},{l_{t}}}\right)}^{F}\left({{\bf{P}}_{{l_{t}}}^{F}-{\bf{P}}_{{l_{s}}}^{F}}\right)\log\left({\frac{{{\bf{P}}_{{l_{t}}}^{F}}}{{{\bf{P}}_{{l_{s}}}^{F}}}}\right)\right),$ (10)
@@
3.4 Training procedure
@@
1 for $k\leftarrow 1$ to $q$ do /* Iterate through all correlated sets. */
@@ 4  ${{\cal L}_{Intra}}=0$;
5
6 for $i\leftarrow 1$ to $m$ do /* Iterate through all teacher-student layers in the current set. */
@@ 8        ${{\cal L}_{Intra}}+={\rm{TF\_Calibration}}\left({{\bf{F}}_{s}^{i},{\bf{F}}_{t}^{j}}\right)$;
9        if $i==m$ then /* Ensure the uniqueness. */
@@ 25        ${\bf{T}}_{s}^{j},{\bf{R}}_{s}^{j}={\rm{Fusion}}\left({{\bf{F}}_{s}^{j},{\bf{R}}_{s}^{j-1}}\right)$;
 /* The final output is chosen as the representative fused feature. */
@@ 31  ${\bf{u}}_{t}^{k}={\bf{T}}_{t}^{n}$;
35
36for $i\leftarrow 1$ to $q$ do /* Iterate through fused features. */
@@ 38     ${{\cal L}_{Inter}}+={\rm{TF\_Calibration}}\left({{\bf{u}}_{s}^{i},{\bf{u}}_{t}^{j}}\right)$;
41
42 end for
Algorithm 1 Intra-inter distillation processing
In this section, we elaborate on the entire training procedure for intra-inter distillation. The overall workflow is outlined in Algorithm 1. First, we pretrain the teacher network using the multi-resolution short-time Fourier transform (MRSTFT) loss as the backbone loss. The teacher is then frozen to serve as a guide for the student, and KD proceeds concurrently with the student’s training. Intra-set distillation is applied to the encoder, decoder, and F-T processing modules. Within each correlated set, besides calculating the intra-set time-frequency calibration distillation loss ${\cal L}_{Intra}$}, we generate representative teacher-student fused features for the current set through layer-wise recursive computation. Inter-set distillation loss ${\cal L}_{Inter}$} is derived via cross-layer interactions within the fused feature set. The total loss for training the student model is formulated as:
${{\cal L}_{stu}}={{\cal L}_{MRSTFT}}+\sum\limits_{k=1}^{q}{\sum\limits_{i=1}^{m}{\sum\limits_{j=1}^{n}{{{\cal L}_{Intra}}\left({{\bf{F}}_{s}^{i},{\bf{F}}_{t}^{j}}\right)}}}+\sum\limits_{i=1}^{q}{\sum\limits_{j=1}^{q}{{{\cal L}_{Inter}}\left({{\bf{u}}_{s}^{i},{\bf{u}}_{t}^{j}}\right)}},$ (11)
4 Experimental settings
4.1 Dataset setup
@@
4.2 Implementation details
Table 1: Baseline model hyperparameter settings.
Layer
Teacher Student
@@ 4.2 Implementation details
Dilated Block
[…]
(2, 3), (1, 1), 128
$p$: [1, 2, 4, 8]
(128, $T$, $F/4$)
(128, $T$, $F/4$)
$(2,3)$, $(1,1)$, 64
$p$: [1, 2, 4, 8]
(64, $T$, $F/4$)
(64, $T$, $F/4$)
[…]
Dilated Block
[…]
(2, 3), (1, 1), 128
$p$: [1, 2, 4, 8]
(128, $T$, $F/4$)
(128, $T$, $F/4$)
$(2,3)$, $(1,1)$, 64
$p$: [1, 2, 4, 8]
(64, $T$, $F/4$)
(64, $T$, $F/4$)
@@
Table 2: Distillation module hyperparameter settings.
@@
Layer
Time Flow Frequency Flow
@@
Linear TfactorT*factor units ($B$, $T$, $T$) ($B$, $T$, TfactorT*factor) BfactorB*factor units ($T$, $B$, $B$) ($T$, $B$, BfactorB*factor)
ReLU - ($B$, $T$, TfactorT*factor) ($B$, $T$, TfactorT*factor) - ($T$, $B$, BfactorB*factor) ($T$, $B$, BfactorB*factor)
Linear $T$ units ($B$, $T$, TfactorT*factor) ($B$, $T$, $T$) $B$ units ($T$, $B$, BfactorB*factor) ($T$, $B$, $B$)
$l^{2}$}Norm - ($B$, $T$, $T$) ($B$, $T$, $T$) - ($T$, $B$, $B$) ($T$, $B$, $B$)
@@
Layer
Teacher Student
@@
In this paper, we use the DPDCRN model as our baseline, which takes complex spectrograms as input and outputs complex ratio masks. For the single-channel DNS dataset, the input channels are set to 2, while for the multi-channel L3DAS23 dataset, the input channels are set to 16. The hyperparameters of the teacher and student models are shown in Table 1, where $C$ represents the number of input channels, $T$ is the number of speech frames, and $F$ is the frequency dimension of speech features. For the encoder and decoder parts, the number of channels in each convolutional layer of the teacher is twice that of the student. In Table 1, $p$ represents the dilation factor, and both the teacher and student models use four types of dilation receptive fields. For the intermediate part, the teacher model employs four F-T modules, while the student model uses only one F-T module, with the number of GRU hidden units also being half those of the teacher model’s.
The hyperparameter settings for the trainable layers in the proposed distillation components are shown in Table 2. The trainable components contain the computation of query and key embedding matrices in time-frequency calibration, as well as the generation process of recursively fused features for each correlated set. In the time-frequency calibration, the similarity embedding mappings are computed separately from two flow directions, with each flow containing two fully-connected layers for dimension transformation. The dimension transformation coefficient factorfactor in Table 2 is set to four in this paper. Residual fusion integrates the current layer features with the recursive features inherited from the previous layer via trainable parameters. In Table 2, $C_{t/s}^{j}$ represents the number of teacher-student channels in the current layer, $F_{t/s}^{j}$ denotes the feature dimension of teacher-student features in the current layer, and $F_{t/s}^{j-1}$ denotes the feature dimension in the previous layer. The number of channels for recursive features of the teacher and student are set to 128 and 64, respectively. Notably, the trainable parameters of the distillation components are only involved in the training of the student model and do not impose additional burden during inference.
All speech signals are sampled at 16 kHz and segmented into 2.5-second chunks. For all models, the window length is set to 32 ms with a frame shift of 16 ms, and the FFT size is 512 points. We employ the MRSTFT loss [6] as the backbone loss. The models are implemented using PyTorch. The training configuration includes a learning rate of 0.0006 with the Adam optimizer, a batch size of 8, and a total training duration of 20 epochs.
4.3 Comparative methods
@@
4.4 Evaluation metrics
In this paper, we use six objective speech quality evaluation metrics, including the perceptual evaluation of speech quality (PESQ) [^3] The evaluation codes are available at https://ecs.utdallas.edu/loizou/speech/software.htm., short-time objective intelligibility (STOI) [^4] The evaluation codes are available at http://www.ceestaal.nl/code/., scale-invariant signal-to-noise ratio (SI-SNR)3, and three mean opinion ratings on quality scales3: signal distortion (CSIG), the intrusiveness of background noise (CBAK), and overall quality (COVL). For all these metrics, higher scores indicate better performance. Additionally, on the L3DAS23 dataset, we use the official recommended word error rate (WER) [^5] The evaluation codes are available at https://github.com/l3das/L3DAS22. to evaluate SE effects for speech recognition purposes, where lower values indicate better performance.
5 Experimental results and discussion
5.1 Ablation study
We conduct ablation experiments for the distillation components on both the validation set and the multi-SNR test set to verify the effectiveness of the proposed modules step-by-step. First, we set the M1 (layer-wise KD) method as the basic form of feature distillation, adopting the teacher-student layer-wise distillation framework of FitNet [26]. Then, M2 (intra-set matching with SemCKD [32]) is constructed on the basis of an intra-set cross-layer distillation framework, where multi-layer matching distillation was performed within each correlated set and calibrated by the cross-layer attention mechanism introduced in SemCKD [32]. M3 (intra-set matching with TFCKD) further extended M2 by replacing the attention-based alignment with the proposed time–frequency cross-calibration described in Section 3.3 of this paper. Finally, the M4 method (i. e., the proposed I2RF-TFCKD framework) performs time–frequency cross-calibrated multi-layer distillation both within and across correlated sets through recursive feature fusion, forming an intra–inter representation fusion distillation framework. The training trends of backbone loss and the speech quality evaluation metric PESQ on the validation set are illustrated in Fig. 5(a) and Fig. 5(b), respectively. Table 3 presents the comparison of objective speech metrics across different distillation strategies on the multi-SNR test set, and statistical significance analysis is conducted in the average metric comparison under three SNRs. We use the significance level $p$ to measure the statistical improvement of each distillation method compared to the student model.
@@ 5.1 Ablation study
Table 3: Speech evaluation metrics for ablation comparisons of distillation components under different SNRs.
Methods
Feature
distillation
Cross-layer
distillation
T-F
calibration
Intra-inter
interactions
Metric
@@ Methods
-M1 (layer-wise KD) - - - 1.966 3.377 2.840 2.657 0.870 0.844
-M2 (intra-set matching with SemCKD [32]) - - - 2.003 3.437 2.884 2.708 0.874 11.020
-M3 (intra-set matching with TFCKD) - - 2.057 3.500 2.894 2.767 0.880 11.107
-M4 (I2RF-TFCKD) - 2.106 3.567 2.950 2.830 0.886 11.253
@@
-M1 (layer-wise KD) - - - 2.329 3.772 3.224 3.059 0.920 14.106
-M2 (intra-set matching with SemCKD [32]) - - - 2.368 3.826 3.262 3.107 0.924 14.237
-M3 (intra-set matching with TFCKD) - - 2.426 3.904 3.274 3.174 0.928 14.270
-M4 (I2RF-TFCKD) - 2.491 3.954 3.333 3.238 0.931 14.461
@@
-M1 (Layer-wise KD) - - - 2.700 4.147 3.591 3.452 0.952 17.041
-M2 (Intra-set matching with SemCKD [32]) - - - 2.724 4.180 3.617 3.483 0.953 17.149
-M3 (Intra-set matching with TFCKD) - - 2.788 4.244 3.633 3.546 0.956 17.157
-M4 (I2RF-TFCKD) - 2.850 4.285 3.686 3.606 0.958 17.356
@@
-M1 (Layer-wise KD) - - - 2.332 ($\ast$)1 3.765 ($\ast$) 3.218 ($\ast$) 3.056 ($\ast$) 0.914 ($\ast$) 13.997 ($\ast$)
-M2 Intra-set matching with SemCKD [32] - - - 2.365 ($\ast$) 3.815 ($\ast\,\ast$) 3.254 ($\ast$) 3.099 ($\ast$) 0.919 ($\ast\,\ast$) 14.135 ($\ast$)
-M3 (Intra-set matching with TFCKD) - - 2.424 ($\ast\,\ast$) 3.883 ($\ast\ast\ast$) 3.267 ($\ast\,\ast$) 3.162 ($\ast\ast\ast$) 0.921 ($\ast\ast\ast$) 14.178 ($\ast$)
-M4 (I2RF-TFCKD) - 2.482 ($\ast\ast\ast$) 3.935 ($\ast\ast\ast$) 3.323 ($\ast\ast\ast$) 3.225 ($\ast\ast\ast$) 0.925 ($\ast\ast\ast$) 14.357 ($\ast$)
  • 1
    The significance analysis [7] of each distillation method and the baseline DPDCRN-S under average SNR is presented by the symbol $\ast$, and the significance level $p$ is: $\ast$: $p>0.1$, $\ast\,\ast$: $0.05<p<0.1$, $\ast\ast\ast$ : $p<0.05$.
5.1.1 The impact of intra-set cross-layer distillation
Compared with layer-wise KD M1, M2 further introduces an intra-set cross-layer distillation mechanism that enables the student to absorb multi-level hierarchical knowledge from the teacher. As shown in Fig. 5(a), M1 enables the student’s backbone loss to decrease rapidly in the early training stages, but this initial advantage does not persist until model convergence. In contrast, M2 achieves a lower converged backbone loss, indicating that cross-layer interactions provide additional knowledge to sustain the student’s learning. In the PESQ curves of Fig. 5(b), M2 consistently outperforms M1, which further validates this observation. In terms of the metric comparisons on the multi-SNR test set in Table 3, M1 is inferior to other methods, showing no significant improvement ($p>0.1$) over the student model in any of the six metrics. By contrast, M2 that extends distillation to multi-layer teacher–student interactions, outperforms M1 across all metrics and achieves secondarily significant improvements ($0.05<p<0.1$) in CSIG and STOI compared to the student model. These results demonstrate that the additional representational information provided by cross-layer interactions offers constructive guidance to the student in both speech quality and noise suppression.
5.1.2 The impact of time–frequency cross-calibration
Building upon the cross-layer KD framework of M2, M3 introduces the proposed time–frequency cross-calibration mechanism, which adjusts the distillation weights across layers tailored to the characteristics of speech signals. As shown in Fig. 5(a), the backbone loss of M3 exhibits a more uniform trend and achieves better convergence compared with M2, indicating that assigning multi-layer distillation weights based on time-frequency calibration leads to more stable knowledge transfer. Similarly, M3 also achieves a better performance ceiling on the PESQ curve of the validation set. Regarding the objective metrics on multi-SNR test set, M3 achieves statistically significant improvements over the student model on CSIG, COVL, and STOI ($p<0.05$), and consistently outperforms both M1 and M2 across all three SNR conditions. These observations demonstrate that cross-calibration captures discriminative information across frequency bands and temporal segments at different layers, enabling more refined teacher-student layer matching.
5.1.3 The impact of the intra-inter representation fusion framework
Distinct from the previous three methods, M4 further introduces a global knowledge circulation mechanism across both intra-set and inter-set levels. This framework not only enables multi-layer teacher–student feature interactions within a single correlated set, but also facilitates cross-set deep fusion of teacher and student representations. As shown in the training curves of Fig. 5, M4 achieves the lowest convergence point of backbone loss among all schemes, while its PESQ curve consistently remains at the highest level, demonstrating stable and sustained performance gains. This indicates that intra-set and inter-set distillation complement each other, expanding the depth of knowledge perceived by the student model. Considering the objective metrics comparisons in Table 3, M4 achieves statistically significant improvements ($p<0.05$) over the student model on five metrics under average SNR, except for SISNR, and shows clear advantages over the other methods across all SNR levels. In summary, the proposed M4 method maximizes the transferability of the teacher by ensuring the student’s absorption of intra-set fine-grained knowledge while simultaneously broadening its understanding of the teacher’s inter-set global information.
5.1.4 Heatmap visualization analysis of time-frequency calibration weights
5.2 Comparative experiments on the DNS benchmark
Fair comparisons of various algorithms are conducted on the DNS non-reverb test set. Table 4 presents the objective speech quality evaluation metrics of various SE methods. We provide details on the causality of each model and whether it uses future frame information (Casual and No Future Information, Casual & NFI). Unlike the 40 ms look-ahead constraint of the DNS challenge, we impose strict restrictions on both the causal architecture and the exclusion of future information to better adapt to real-time applications. Additionally, the model parameter (Param.) count and floating-point operations (FLOPs) per second are provided to evaluate deployment feasibility on edge devices. Among the SE models in Table 4, NSNet, RNNoise, and DTLN are low-complexity real-time SE algorithms primarily based on LSTM networks. These models are fully causal and do not utilize future information. In contrast, DCCRN, FullSubNet+, FRCRN, CTS-Net, and PrimeK-Net fail to satisfy both causality and NFI constraints. Although some architectures, such as DCCRN and CTS-Net, employ causal designs, they still leverage partial future frame information during training and inference. Our distillation baseline DPDCRN, composed of causal convolutions and GRU layers, is strictly causal and operates without any future frame information.
5.2.1 Analysis of distillation frameworks
Figure 7: The scatter-bubble distribution of three metrics (PESQ, STOI, and SISNR) with model parameters and FLOPs for SE models on the DNS test set.
We compare the effects of different distillation frameworks applied to the base model DPDCRN in Table 4. Among them, MSE that directly narrows the output distance, only achieves slightly higher results than the student model on PESQ and STOI metrics and even falls below the student on the SISNR metric. In contrast, M1 achieves effective improvements compared to output-level distillation through layer-wise feature distillation. On this basis, both CLSKD and M2 incorporate multi-layer knowledge integration to enable cross-layer interaction between the teacher and the student, thus achieving more considerable gains in PESQ and STOI. The distillation advantages of the above multi-layer knowledge transfer frameworks demonstrate that the intermediate representations of the model contain rich transferable information. Recent studies such as UCLFWPKD, ABC-KD, and Two-Step KD design a targeted cross-layer distillation framework according to the characteristics of the SE task, achieving further improvements. In this paper, M3 introduces time-frequency cross-calibration based on the multi-layer distillation framework of M2, achieving notable improvements across all three metrics. This highlights that computing multi-layer calibration weights along both the temporal and spectral domains enables more effective knowledge allocation. Extending M3 to the intra-inter set distillation framework (M4, I2RF-TFCKD) achieves the best performance among all distillation methods, demonstrating that knowledge circulation within and between correlated sets further facilitate teacher-student knowledge transfer.
5.2.2 Comparison with state-of-the-art SE methods
Table 4: Comparison of objective speech evaluation metrics between distilled and undistilled models on the DNS test set.
Methods Casual & NFI Param. (M) FLOPs (G/s) PESQ STOI SISNR (dB)
Noisy - - - 1.58 0.915 9.07